Besides running his own businesses, Tommy is also an avid stock investor and has a team of professional investment managers overseeing his stock portfolio. He’s trying to determine the fair value of a stock using a simple dividend discount model. Information on the stock is as follows:
Macchiato Ltd is a famous coffee retail chain, and its stock traded at $42 at the beginning of 2016. Its beta estimate by a beta service company is 1 and its market risk premium is 6%. The risk free rate at the end of 2015 was 1.6%. The firm was expected to pay dividends of $1.60 per share at the end of 2016 and 2017.

(b) At what price do you expect Macchiato Ltd to sell at the end of 2017 if you forecast it will not pay the dividends as expected? (3 marks)

(c) At what price do you expect Macchiato Ltd to sell at the end of 2017 if you forecast it will pay the dividends as expected? (3 marks)

(d) Assume that Macchiato Ltd’s investors expect it to pay a dividend of $2.50 per share forever. Using the required rate of return calculated in (a) above, determine the value gained or lost per share by buying a share at $42.
Answer #1

1.
=42*(1+1.6%+1*6%)^2=48.626592

2.
=42*(1+1.6%+1*6%)^2-1.6*(1+1.6%+1*6%)-1.6=45.304992

3.
Price should be=2.50/(1.6%+1*6%)=32.89473684

Value lost=42-32.89473684=9.10526316